OpenStudy (anonymous):

(3 square root of 3)(7 square root of 6)

OpenStudy (anonymous):

\[21\sqrt{18} \] let me check something out though

OpenStudy (anonymous):

okay.

OpenStudy (saifoo.khan):

\[63\sqrt{2}\]

OpenStudy (anonymous):

yea saifoo is right i didn't simplfy the there \[21\sqrt{18}=63\sqrt{2}\]

OpenStudy (saifoo.khan):

=)